an electric bulb is rated 220V and 100W .When it is operated on 110V ,the power consumed will be?

Grade:Select Grade

1 Answers

KAPIL MANDAL
9 years ago
At first, you have to understand why bulb glows. It is because heat produced there due to resistance. 
The voltage V is related to power P by the formula,
P = I * V
I is the current through the resistance where the voltage is applied. Again,
V=IR.
So, P= V2 / R
Or,  R= V2/P 
Here, R = 2202 /100 = 484 ohms.
Putting this back in the eqn P= V2 / R, with V=  110, 
P= 1102/ 484 = 25 watts.